OpenStudy (anonymous):

Calculus. The tangent to the curve y=ax^4 at the point (3,b) has a gradient of 2. Find the values of a and b. Please show working. :)

OpenStudy (ash2326):

(3, b) lies on the curve \(y=ax^4\) so \[b=a\times3^4..................(1)\] It's given the gradient of the tangent is 2 at (3, b), that is \[\large (\frac{dy}{dx})_{(3, b)}=2\] Find dy/dx , put x=3 and y=b, and equate it to 2 you'll get the second equation. Two equations , two variables voila!

OpenStudy (ash2326):

Do let me know if you have any doubt:)

OpenStudy (ash2326):

@Lachlan1996 Do you get this?

OpenStudy (anonymous):

Umm somewhat, do excuse me, i am a little slow with this. we are only just into year 11 calculus. I do follow somewhat but im not sure what you mean by equating (dydx)(3,b)=2

OpenStudy (ash2326):

Okay, we'll do it step by step firstly Find dy/dx

OpenStudy (anonymous):

thats the bit im not sure with

OpenStudy (ash2326):

Do you know differentiation?

OpenStudy (anonymous):

i am familiar with the basics, but we are still learning it

OpenStudy (ash2326):

@Lachlan1996 I gotta bounce now, Sorry My friend @apoorvk will help you on this

OpenStudy (anonymous):

okie doke, mate, thanks for the help so far

OpenStudy (apoorvk):

Hmm... okay so do you want me start all over again? I can do that, not a problem..

OpenStudy (anonymous):

Do whatever suits you mate, im happy with any help

OpenStudy (apoorvk):

Okay, I'll proceed a wee bit of recap. So, did you get the part where we satisfy (3,b) in the equation of the curve \(y=ax^4\), since the point lies on the curve?

OpenStudy (anonymous):

yes. The x coordinate 3 is substituted into the formula [y=ax^4] to get a*81

OpenStudy (apoorvk):

So, b = 81a ....(i) This is one equation that we have. Okay good. Now now.. We need a general equation of the slope.gradient value for the curve. What do we do - shoul I differentiate \(y=ax^4\)??

OpenStudy (apoorvk):

*should

OpenStudy (anonymous):

uhh \[y=ax^4\] = \[4ax^3\] doesn't it?

OpenStudy (apoorvk):

yeah, means:\[y' = \frac{d y}{dx}= 4ax^3\] Now, dy/dx, that is the " \(4ax^2 \) " obtained, is the general expression for the value of slope/gradient for the curve \(ax^4\) at any particular 'x'. Okay till here?

OpenStudy (anonymous):

do you mean "ax^3

OpenStudy (apoorvk):

oh yeah, I meant \(4ax^3\). Typo sorry.

OpenStudy (anonymous):

and so 4(a)(3)^3-1 = the gradient of this point

OpenStudy (anonymous):

no sorry typo there too

OpenStudy (anonymous):

and so 4(a)(3)^4-1

OpenStudy (apoorvk):

Umm. you mean \(4a(3)^{4-1}\) right? That's the slope at (3,b)

OpenStudy (anonymous):

yes thats what i meant

OpenStudy (anonymous):

and so 108a=2. thereby a=1/54

OpenStudy (anonymous):

and so b = 1/54 *81

OpenStudy (apoorvk):

Hmm, and now, according to the question the value of the gradient at (3,b) is '2'. So, \(4a3^3 = 108a = 2\) And yes, so a = 1/54

OpenStudy (anonymous):

Thankyou very much for your help :D

OpenStudy (apoorvk):

And we had an equation in the beginning remember? b= 81a so, b = 81/54 = 3/2 That's it! You are very welcome, and you're a good learner!

OpenStudy (anonymous):

your a top bloke mate :) cheers for the shunt i needed. Helped me out a lot. Big thanks to you!

OpenStudy (apoorvk):

hey man, not me, it was your own desire to learn. Cheerio ;)

OpenStudy (anonymous):

Have a good night mate :) wish you well

OpenStudy (apoorvk):

Good night to you as well! Let us know if we can help you anymore! :)

OpenStudy (anonymous):

Will do mate. Gotta knock the maths off, then chem and physics

OpenStudy (apoorvk):

Kick 'em hard! Good luck!

OpenStudy (anonymous):

Will do. Big thanks to you, and give my thanks to ash

OpenStudy (apoorvk):

sure mate, it was his work on this after all ;)

OpenStudy (anonymous):

The tangent to the curve y=ax^3 at the point (-1,b) is perpendicular to the line y=2x+3 Find the values of a and b. I would take this in the same fashion by finding the gradient first yes?